### v1

# Star formation beyond galaxies: widespread in-situ formation of intra-cluster stars in TNG50

Abstract
We study the fraction of the intra-cluster light (ICL) formed in-situ in the three most massive clusters of the TNG50 simulation, with virial masses . We find that a significant fraction of ICL stars (-) are born in-situ. This amounts to a total stellar mass comparable to the central galaxy itself. Contrary to simple expectations, only a sub-dominant fraction of these in-situ ICL stars are born in the central regions and later re-distributed to more energetic orbits during mergers. Instead, many in-situ ICL stars form directly hundreds of kiloparsecs away from the central galaxy, in clouds condensing out of the circum-cluster medium. The simulations predict a present-date diffuse star formation rate of 1 /yr, with higher rates at higher redshifts. The diffuse star forming component of the ICL is filamentary in nature, extends for hundreds of kiloparsecs and traces the distribution of neutral gas in the cluster host halo. We discuss briefly how numerical details of the baryonic treatment in the simulation may play a role in this result and conclude that a sensitivity of erg s cm arcsec in H flux (beyond current observational capabilities) would be necessary to detect this diffuse star-forming component in galaxy clusters.
keywords:
galaxies: clusters: intracluster medium – Galaxy: formation – Galaxy: evolution1 Introduction
The simplest models of galaxy formation suggest that most stars in a cluster should form in the high-density regions of individual galaxies. These models imply that the majority of stars in the intra-cluster light (ICL) originate from the tidal removal of stars from luminous galaxies with mass similar to the Milky Way (MW) (Puchwein et al., 2010; Cui et al., 2014; Cooper et al., 2015; Contini et al., 2013, 2018, 2019; Montenegro-Taborda et al., 2023), with relatively small contributions from dwarf mass galaxies, particularly in group environments with halo virial masses (Ahvazi et al., 2023). Nevertheless, observational studies (Barfety et al., 2022; Hlavacek-Larrondo et al., 2020) and some numerical simulations (Puchwein et al., 2010) indicate that a non-negligible fraction of ICL stars are formed in-situ.
In-situ ICL stars could have formed in the brightest cluster galaxy (BCG) galaxy and been later pushed onto highly energetic orbits by mergers. An alternative, more intriguing, possibility is that stars form directly in the circum-cluster medium (CCM), far away from the high density regions of the BCG. This low-density mode of star formation is potentially interesting, as it might occur under very different physical conditions than the normal star formation in galaxies, with consequences for the multi-phase nature of the CCM and its metallicity, as well as for the age distribution of ICL stars, especially if in-situ star formation continues to the present day. From an observational perspective, if a substantial amount of stars form outside galaxies, this would have strong implications for wide-field surveys in the UV or H, which are especially sensitive to the presence of young stars.
Several mechanisms have been suggested to explain in-situ ICL star formation. One of the most commonly cited is the compression of gas associated with the ram-pressure trails of infalling galaxies, a phenomenon that has been detected in both observations (Smith et al., 2010; Mihos et al., 2016; Boselli et al., 2018a; Gullieuszik et al., 2020) and some numerical simulations (Kronberger et al. 2008; Kapferer et al. 2009; Tonnesen & Bryan 2012; Lee et al. 2020; Calura et al. 2020; Steyrleithner et al. 2020; Vollmer et al. 2021; Göller et al. 2023, for a detailed discussion see section 5.3.1 in review paper by Boselli et al. 2022). Another possible mechanism is star formation within the tails of stripped material formed during gravitational fly-by encounters of gas-rich systems (see see NGC 4254 in Virgo, Boselli et al. 2018b; VCC 1249, Battaia et al. 2012; IC3418, Hester et al. 2010; Fumagalli et al. 2011; Kenney et al. 2014).
In-situ star formation in the CCM is also predicted to be widespread at high-redshift, where AGN feedback might be less powerful and unable to offset the short timescales for cooling in the dense regions of the CCM (Webb et al., 2015b). In such a scenario, spontaneous run-away cooling in the CCM might lead to an increase of density high enough to trigger violent star-formation events in locations outside galaxies (Hlavacek-Larrondo et al., 2020).
Interestingly, observations of off-center star formation at large distances from the BCG are consistent with this scenario for a few high- clusters (Webb et al., 2015a; Hlavacek-Larrondo et al., 2020; Barfety et al., 2022), in some cases predicting as much as a MW-worth of stars contributed to the CCM by such cooling events.
Some observations of isolated HII regions in nearby clusters, such as the Virgo cluster, may also hint at a more sustained but lower-level star formation activity taking place directly in the CCM/ICL region (Gerhard et al., 2002; Bellazzini et al., 2018). However, it is not always trivial to disentangle whether these isolated star-forming clouds are CCM-born material or belong to some older ram-pressure or interaction events in the cluster (Junais et al., 2021; Jones et al., 2022).
Numerical simulations therefore emerge as valuable tools to explore this mode of star formation within the realistic framework of the CDM model. For instance, Puchwein et al. (2010) estimate that approximately of the ICL stars formed in their clusters (with masses between and ) originated at distances consistent with being CCM-born.
Various numerical methods have been employed to study the formation of cold gas in the CCM, which could give rise to star formation outside galaxies. Runaway local production of cold gas in the CCM, as well as thermal instability, have been shown to lead to multi-phase structures in idealized hot haloes (Maller & Bullock, 2004; Kaufmann et al., 2006, 2009; McCourt et al., 2012; Sharma et al., 2012).
Given that the cold-phase CCM resulting from such processes is a small-scale phenomenon, it has been studied largely in idealized numerical experiments (Fielding et al., 2020). Previous work by Oppenheimer et al. (2018) demonstrated a relatively significant abundance of cold, low-ionization gas in the CCM of EAGLE galaxies, albeit with uncertainties regarding the numerical robustness of the hydrodynamical technique in this regime (see also Ford et al. 2013).
Semi-analytic models have also been developed to explain the unique, possibly drag-induced sub-virial kinematics of cold gas flowing through luminous red galaxy host haloes (Afruni et al., 2019). In the context of IllustrisTNG, Nelson et al. (2018) showed that TNG is consistent with empirical OVI constraints in the CCM.
The cold gas content of galaxies themselves (both neutral HI and molecular H2) has also been favorably compared against data in several regimes. Nelson et al. (2020) also demonstrated that group mass haloes hosting massive galaxies are filled with a hot, virialized plasma at K, and that they contain a large abundance of cold gas with a characteristic temperature of K. This cold phase of the CCM takes the form of thousands of small, kpc-sized clouds, which fill the inner halo out to hundreds of kpc (similar to predictions by Maller & Bullock 2004 for MW-mass halos).
There are, however, several important caveats when studying the formation of star-forming clouds in the CCM, mostly connected with limitations of numerical techniques. Approaches such as SPH have been shown to be unable to track the evolution of instabilities, such as the Kelvin-Helmholtz instability, leading to the formation of more dense clouds that are artificially long-lived compared to similar numerical experiments run with grid-base techniques (Agertz et al., 2009; Torrey et al., 2012).
The lack of, or inadequate treatment of other important physics, such as heat conduction, magnetic fields or metal diffusion might also complicate the interpretation of results from simulations (see the discussion in Puchwein et al. 2010). It is clear that, for any numerical technique chosen, high resolution becomes key to study in-situ star formation in low-mass clouds embedded in the hot CCM of clusters, especially considering the low masses of these clouds.
In this context, we present here a study of the in-situ formation of stars and their contribution to the ICL using the three most massive groups/clusters in the TNG50 simulations (Pillepich et al., 2018b; Nelson et al., 2019), one of the highest-resolution cosmological simulations of low-mass clusters with virial mass at .
This article is organized as follows: in Section 2, we provide a comprehensive overview of the simulation employed, including some definitions crucial for the subsequent sections. Section 3 presents our findings, together with an in-depth discussion. Finally, Section 4 summarizes the key conclusions drawn from this study.
2 Simulation
We use the cosmological hydrodynamical simulation IllustrisTNG (Pillepich et al., 2018b; Springel et al., 2018; Marinacci et al., 2018; Naiman et al., 2018; Nelson et al., 2019). The IllustrisTNG project111https://www.tng-project.org is a suite of state-of-the-art cosmological galaxy formation simulations that come in different physical sizes and mass resolutions. In this work we use the highest resolution baryonic run TNG50-1 (TNG50 for short), which has a volume of approximately . It has resolution elements, and the baryon and dark matter particle masses are and , respectively. The IllustrisTNG cosmological parameters are consistent with CDM model determined by Planck XIII (Planck Collaboration et al., 2016) to be with and
Gravity and baryons are followed using the Arepo moving mesh code (Springel, 2010). The baryonic treatment is based on the code used for its predecessor simulation suite, Illustris (Vogelsberger et al., 2013a, 2014a, 2014b; Genel et al., 2014; Nelson et al., 2015), with modifications implemented to better track the formation and evolution of galaxies (as described in Pillepich et al., 2018a).
The updated IllustrisTNG sub-grid models account for star formation, radiative metal cooling, chemical enrichment from SNII, SNIa, and AGB stars, stellar feedback, and super-massive black hole feedback. The star formation in the Interstellar Medium (ISM) is treated in a sub-resolution fashion following a variation of the Springel & Hernquist (2003) with the adoption of a Chabrier (2003) initial mass function, and model the star-forming dense ISM gas using an effective equation of state where stars form stochastically above a gas density of = with a star formation time-scale of = Gyr (refer to Vogelsberger et al. 2013b for detailed explanations). These models were calibrated to reproduce a set of observational constraints that include the observed galaxy stellar mass function (Vogelsberger et al., 2014a), the cosmic star formation rate (SFR) density (Genel et al., 2014), colors and morphologies of the present day galaxy population (Sales et al., 2014; Snyder et al., 2015; Rodriguez-Gomez et al., 2017).
Galaxies in the simulations are identified using SUBFIND (Springel et al., 2001; Dolag et al., 2009) on Friends-of-Friends (FoF) groups (Davis et al., 1985) identified with a linking length set to 0.2 times the mean interparticle separation. The object at the center of the gravitational potential of each group is called the “central” galaxy, while all other substructures are “satellites”. The time evolution of galaxies and halos through the 99 snapshots of the simulation is followed using the SubLink merger tree algorithm (Rodriguez-Gomez et al., 2015).
| Group 0 | Group 1 | Group 2 | |
| M [] | |||
| M [] | |||
| M [] | |||
| M [] | |||
| f | |||
| F |
2.1 Defining ICL and cleaning method
We focus the analysis on the most massive systems included in the TNG50 box, which fulfill , comparable to nearby systems such as the Fornax and the Virgo clusters. Virial masses for these systems are listed in table 1, where virial quantities are defined at the virial radius , corresponding to the distance at which the enclosed density is times the critical value.
To define the particles that are considered part of the ICL we proceed as follows. For each of our groups, we remove all stars that SUBFIND considers bound to surviving satellite galaxies. The left panel of Fig. 1 shows an illustration of this component for our Group 1. Furthermore, to avoid possible contamination by stars in the outskirts of the surviving satellites that are wrongly assigned to the host halo by SUBFIND, we further apply a “cleaning” of our ICL: we remove any stellar particle within of the center of any subhalo with , where refers to the stellar half mass radius of the subhalo. This ensures that any stars from satellites and their surrounding stellar halo are not considered part of the ICL.
This leaves us with a sample consisting of all the remaining stellar particles in the FoF group that are beyond of the central galaxy and within the virial radius of the host cluster, which we call ICL.
Stars in the ICL are further divided into those born “in-situ” and those “accreted”. Following Rodriguez-Gomez et al. (2016), our algorithm checks if the star is bound to the main progenitor of the central galaxy at the time of birth (or the closest snapshot to it), in which case it is labeled as in-situ, or it is otherwise associated by SUBFIND to a different substructure in the SubLink merger tree, in which case it is labeled as accreted. The catalogs for in-situ/accreted are provided within the IllustrisTNG database. As an illustration, the middle and right panels of Fig. 1 show the accreted and in-situ components of the ICL in Group 1, respectively.
To quantify the mass within the ICL and, specifically, the in-situ stellar component, we introduce the fractions and . Various definitions of these fractions exist in the literature, and for the purpose of our study, we adopt the following definitions:
and
where refers to the mass of the stellar component in the ICL, represents the mass of the in-situ stellar component in the ICL, and is the stellar mass within of the central galaxy. The values for these quantities in the three most massive systems in TNG50 are presented in table 1.
2.2 H predictions
In order to compare results with the recent observations of the ICL in nearby clusters (i.e. Virgo), we use the calibration of Kennicutt & Evans (2012) to convert the star formation rates to H luminosities. These calibrations are based on evolutionary population synthesis models, in which the emergent spectral energy distributions (SEDs) are derived for synthetic stellar populations with a prescribed age mix, chemical composition, and Kroupa (2001) initial mass function (IMF). The resulting relation is
where denotes the luminosity in units of , represents the star formation rate in units of , and is the logarithmic SFR calibration constant which we set to be 41.27 (Kennicutt & Evans, 2012; Murphy et al., 2011; Hao et al., 2011).
3 Results
We begin by characterizing the mass distribution within our groups. Fig. 2 illustrates the projected radial distribution of stars (solid thick black line), dark matter (black dotted line), and satellite galaxies with (depicted as orange circles) after projecting each group along the axis. The stellar component is predicted to fall off more steeply than both the dark matter and the satellite galaxies in these groups (upon vertical re-scaling by a factor of , the satellite galaxies appear to closely trace the dark matter distribution in this simulation). For comparison, the purple dotted line represents the stellar mass surface density profiles from Cooper et al. (2013), averaged for halos with . These results align closely with our stellar profiles (solid black lines) for the two less massive clusters. However, in the case of our most massive cluster, having a halo mass twice that of the most massive halo in their sample, differences with the predictions in Cooper et al. (2013) are evident. Note that the tagging method in Cooper et al. (2013) does not properly account for the in-situ star formation in the BCG, therefore giving a flatter slope in the inner regions. We further divide the stars according to their origin: in-situ (represented by the thin blue dot-dashed curve) and accreted (depicted by the thin red long-dashed curve). While the dominance of the in-situ population is expected within the BCG (the region highlighted by gray shading), their relatively substantial contribution to the ICL at hundreds of kpc is intriguing. In our systems, in-situ born stars contribute between and of the stellar mass in the ICL (refer to the top right corner of each panel for the specific percentage of each group).
An inspection of the birth place of the in-situ labelled stars in the ICL shows that the majority of them are formed already at large distances from the BCG (although not associated with substructures) and are not stars from the central galaxy kicked-off by mergers, a common formation path for in-situ stars in the diffuse stellar halos of MW-like galaxies, and as is the case for our most massive cluster (which experienced a recent major merger in its formation history). However, for the remaining groups in our sample, over 60% of the in-situ ICL is formed at kpc or (where distance is in physical kpc and is measured with respect to the BCG or its main progenitor at the time of birth of the star). This suggests that gas in the CCM must somehow become available for star formation beyond the inner regions dominated by the BCG. This star formation mode is thus happening in unconventionally diffuse conditions, very different from the star formation resulting from the dense interstellar medium of galaxies.
We calculate the total SFR of CCM gas at for our groups and compare this to the level of star formation ongoing in the central BCG as well as in other galaxies. Fig. 3 highlights that, taken as a unit, the ICL (integrated within the virial radius) is forming stars at rates comparable to that of the BCG and other massive galaxies. For this comparison, we assign a stellar mass to the ICL equal to the sum of all in-situ and accreted stellar particles that lay beyond and up to the virial radius and that are not associated to any satellite or identified substructure by SUBFIND. Similarly, we compute its SFR by summing the star formation rates of all gas cells that are not gravitationally bound to any substructure within the same region. Fig. 3 shows that while the ICL lays systematically below the “main sequence” of star forming galaxies (magenta points show central galaxies within the TNG50 box), its SFR is comparable to many of the massive satellite galaxies in the group (gray circles) and slightly above the SFR of the central BCG in the case of Group 1 (middle panel). For comparison, we present the extrapolated “main-sequence” relation derived from the galaxies in the Virgo cluster (see Boselli et al. 2023), for a sample with normal HI gas content (HI-normal) and deficient HI gas content (HI-deficient); the mean values and standard deviations for HI-normal and HI-deficient objects are shown by blue and red markers. We are also presenting results from late-type nearby galaxies in the Herschel Reference Survey by Boselli et al. 2015 (dashed green line), a large statistical sample of SDSS galaxies from Peng et al. 2010 (dashed brown line), and the extrapolated time-dependent main-sequence relation derived by Speagle et al. 2014 (dashed orange line).
The ongoing star formation in the CCM shows a decreasing trend with distance from the center of the cluster. Fig. 4 presents the surface density of star formation rate (), computed by projecting our groups along the -axis and integrating star formation within square bins of kpc on each side. Different colors represent different groups, while individual points of the same color depict the scatter between various regions within a given group. Bins with no star formation at all are not shown. The thick solid curve corresponds to the median trend (considering all groups collectively) of as a function of clustercentric distance, highlighting that the inner regions of the CCM exhibit more prominent star formation compared to the outskirts.
To ease comparison with observations, we convert into a surface density of H luminosity (right vertical axis) as explained in detail in Sec. 2.2 and following Kennicutt & Evans (2012). Typical values in the CCM of our sample span with considerable scatter. These values are lower than the average measured in central regions of disk galaxies, but comparable to the values in the outskirts of MW- and LMC-like objects (Tacchella et al., 2022; Bundy et al., 2015). Beyond clustercentric distances diffuse star formation is rare in all our systems.
Projected maps of our halos in Fig. 5 show in more detail where the star formation in the intracluster space is happening. Panels on the first row show the star-forming gas in the region (bounds shown with magenta and orange circles) color coded by SFR (color map on the right). Star formation at is widespread in these systems and follows filament-like patterns. We highlight that special care was taken to remove any material gravitationally associated with surviving satellites, as described in Sec. 2.1, and this star formation is occurring in gas seemingly associated to the central host halo and not to bound substructures. In fact, further inspection of each system revealed that the filamentary nature of the star-forming gas follows the distribution of neutral gas in the main host halo (bottom row), and looks loosely linked to the dark matter distribution (middle row) for at least 2 of the systems (groups 1 and 2).
To gain further insight, we zoom into smaller regions of group 1 in Fig. 6. The yellow points show the distribution of the stellar particles in each box and reveal the presence of several galaxies, some with associated stellar streams—most prominently observable in the two innermost regions (red and green panels). The presence of the ICL is apparent in panels as a diffuse yellow stellar background on the end closer to the central region of cluster. We highlight the presence of star forming gas colored in magenta by their estimated flux (refer to the color bar on the side of the main panel). For the zoomed-in panels, the color maps are calculated in kpc bins to align with the arcsec spatial resolution of the VESTIGE survey of the Virgo cluster (Boselli et al., 2018a), while the main panel is smoothed to 1 arcmin.
The zoomed-in regions in Fig. 6 show that the majority of the star forming gas is mostly unrelated to the distribution of stars or galaxies. Star forming regions are also not obviously associated to ram-pressure tails from gas-rich galaxies note that this is different from what has been observed in cases such as Smith et al. (2010); Mihos et al. (2016); Boselli et al. (2018a); Gullieuszik et al. (2020), where the association of the star-forming region to tails of ram-pressure stripped galaxies is clear upon visual inspection. Instead, stars seem to be forming in cloudlet-like objects with tens-of-pc typical size, which condensate along filamentary structures in the CCM.
This is not the first time that small-scale cold clouds in the CCM have been reported in TNG50. In a detailed analysis of the CCM in halos, Nelson et al. (2020) highlighted this fundamental prediction from the TNG model in combination with the high resolution achieved by TNG50. These small-scale cold clouds seem consistent with being seeded by local inhomogeneities in gas density that later grow due to increased cooling. While the role of magnetic fields in forming these clouds is not clear, the magnetic pressure in these clouds is large and exceeds the thermal pressure, providing support and making these features long-lived, on the scale of several Gyrs. Interestingly, the results in this work show that some of these cold clouds manage to become star-forming and contribute a non-negligible fraction of the stars to the diffuse ICL component.
In agreement with the findings of Nelson et al. (2020), the metallicity of in-situ born ICL stars (with ) is relatively high, registering a median value of with a substantial scatter of approximately dex, comparable to the accreted ICL, the metallicity of dwarf galaxies with , and the outskirts of more luminous, MW-like galaxies. This suggests that the gas seeding these density perturbations has galaxy-like abundances and it likely comprises gas that has, a long time ago, been stripped from infalling galaxies. Obvious associations to those satellites do not seem to persist until in our simulations, suggesting that the timescales for the clouds to form and become star-forming are comparable to the orbital times of the satellites.
The physical conditions for the star-forming clouds seem comparable to those of normal star-forming regions in the ISM of galaxies. The median density of the star-forming cells flagged as part of the ICL is , which is just above the density threshold for star formation in the TNG baryonic treatment . Interestingly, the fraction of stellar mass born in-situ in the ICL seems quite independent of numerical resolution, as clusters selected in TNG100 show similar levels of star formation rate in the ICL (see Appendix A). The presence of these in-situ born stars in the ICL is therefore not driven by the increase in resolution from TNG50 to TNG100 but is instead a more fundamental consequence (or prediction) of the model.
It is also important to note that at the current integrated star formation rates shown in the intra-cluster region of our simulated clusters, SFR 1 /yr (see Fig. 3), it would take Gyr to form the total mass of the in-situ ICL stars listed in Table 1. This diffuse mode of star formation must therefore have also been present at earlier times, and with higher intensity. We confirm this by showing the distribution of stellar mass formed in-situ in a given time-bin in Fig. 7. This is calculated by adding all the stellar mass of the in situ-labeled stars in the ICL with a given age divided by the size of the time-bin to provide an average star formation rate. The blue shaded histograms indicate that the in-situ intra-cluster star-formation is a rather continuous phenomena over time and was, on average, stronger around - (i.e., cosmic times - Gyr). As an illustration, we show in Fig. 8 the predicted SFR maps for our highest mass group at three different redshifts: . In all cases, extended star forming areas appear out to almost the virial radius of the cluster progenitor.
Have these star-forming regions been detected in observations? In the nearby universe, ongoing star formation in the tails of ram-pressured stripped galaxies has been observed in the past, for example in the Virgo cluster (Mihos et al., 2016; Boselli et al., 2018a). However, the predictions from TNG50 are different in nature from these observations, since the star-forming cloudlets are currently not associated with any satellite. In that sense, they are closer to a recently identified class of young, isolated, and star-forming clouds in Virgo, see Jones et al. (2024); Beccari et al. (2017). We cannot rule out that this gas originated as ram-pressure tails stripped from galaxies in the distant past, where any possible correlation with the galaxy distribution has been long erased. However, if that is the case, the timescale for turning this gas into stars seems too long compared to observations of ram-pressure tails which can still be traced back to their original galaxy. Therefore we consider this a different process than what is observed in ram-pressure tails.
The color coding of the gas in Fig. 6 indicates that, at , the expected H surface brightness for the intra-cluster gas is rather modest, – , which could partially explain why this phenomena has so far escaped detection. For comparison, the most extended and deepest surveys available, such as VESTIGE, have a sensitivity at 3 arcsec angular resolution (Boselli et al., 2018a). On the other hand, there are claims of detection of runaway cooling of the hot gas in high- clusters leading to star formation events with SFR 900 /yr offset from the main host galaxy (Hlavacek-Larrondo et al., 2020) that might be comparable to some of the in-situ formed stars in our simulations.
Finally, we must also consider the possibility that these in-situ formed ICL stars are a numerical effect born out of specific choices made for the baryonic treatment in the simulation. The fact that most star-forming clouds sit near the density threshold for star formation suggests that increasing the value of that parameter might significantly lower the occurrence of stars born in the in-situ ICL component, but more rigorous tests must be performed to quantify this effect.
4 Summary and Conclusions
We use the cosmological hydrodynamical simulation TNG50 to study the formation of in-situ ICL in Virgo-like systems. We select the 3 most massive clusters in the box with virial masses -. We find that a significant fraction, - of the present-day mass in the ICL, was formed in-situ. When analyzed in detail, and contrary to what is commonly believed, the majority of these in-situ stars are born directly into the diffuse light component at hundreds of kiloparsecs from the main host and directly from the CCM gas, rather than brought in and dispersed during merger events. This diffuse star-formation proceeds in cloudlets that condense along filamentary-like structures that are aligned with the neutral gas and dark matter distributions and are not clearly associated with ram-pressure stripping of gas in infalling galaxies.
At the expected H surface brightness due to this diffuse mode of star formation reaches , just below the detection limits of the ongoing surveys in Virgo when smoothed on scales of arcsec. In the simulated clusters, star formation maps at higher redshift suggest that extended star formation is present throughout the history of the cluster, making it an interesting target for discovery with facilities such as JWST.
One interesting application of our results is that orphan supernova events would be expected to occur in relative isolation in the ICL. While such events have not yet been reported, with a predicted average SFR of /yr, we expect only 1 SN per yr in a Virgo-like system222We have assumed a power-law mass function following Salpeter (1955), for the mass range of – with stars more massive than exploding as SN, which results in of the total stellar mass formed able to generate a SN event. Assuming a mean SN mass , we anticipate the occurrence of 1 SN in () years., suggesting that a concurrent survey of hundreds of clusters might be needed for detecting such rare events.
If ongoing star formation in the intra-cluster region of observed clusters is really lacking, then this would signal shortcomings in the treatment of star formation in simulations like TNG50, especially in low-density regions. The formation of stars in these extended and diffuse structures is possible because of the condensation of denser clouds embedded in the more diffuse CCM, as reported by Nelson et al. (2020). Numerical techniques such as that used in Arepo have been shown to lead to less artificial formation of clouds compared to other techniques such as SPH (Agertz et al., 2009; Torrey et al., 2012), but it is possible that other numerical effects are still at play. We emphasize that further studies of the effect of the star formation threshold in models such as those of TNG would be desirable to determine their role in the formation of stars in such unexpected conditions.
Acknowledgements
We gratefully acknowledge the contributions of several individuals who supported this research. NA and LVS would like to acknowledge the financial support received from the NASA ATP-80NSSC20K0566, NSF-CAREER-1945310 and NSF-AST-2107993 grants. NA would like to acknowledge the support provided by the UCR-Carnegie Graduate Fellowship.
Data Availability
This paper is based on halo catalogs and merger trees from the IllustrisTNG Project (Nelson et al., 2019). These data are publicly available at https://www.tng-project.org/.
Appendix A Effect of resolution
To assess the impact of resolution on the star formation rates found within the ICL regions in our studied groups in TNG50, we turn our attention to the results obtained from TNG100. TNG100-1, part of the same suite of cosmological hydrodynamical simulations, represents a larger baryonic run with a volume of and resolution elements (lower resolution compared to TNG50, which features resolution elements). In this run, the baryon and dark matter particle masses are and , respectively.
For the comparison, we exclusively consider groups with virial masses similar to those examined in our study (), resulting in a subset of 26 groups in TNG100.
In Fig. 9, we juxtapose the star formation rates in TNG50 and TNG100 for both central and ICL regions. The results reveal no discernible difference in the ICL star formation rates among various groups between the two simulations (represented by navy squares for TNG50 results and light blue squares for TNG100 results), aside from a larger scatter in the TNG100 results, primarily due to the larger sample size in the more extensive box of TNG100).